Suppose that you start with a normal deck of 52 cards and remove everything except the twos, threes, and fours. So you now have a deck of twelve cards consisting of four twos, four threes, and four fours. A card is drawn at random and replaced then another card is drawn at random and replaced. Make a probability distribution for the sum of the two numbers that are drawn.
Answer by greenestamps(13206)



Since each card that is drawn is replaced before drawing the next card, the probability of each number on each draw is 4/12, or 1/3. Then

P(4) = P(2 and 2) = (1/3)(1/3) = 1/9
P(5) = P(2 and 3 OR 3 and 2) = (1/3)(1/3)+(1/3)(1/3) = 2/9
P(6) = P(2 and 4 OR 3 and 3 OR 4 and 2) = (1/3)(1/3)+(1/3)(1/3)+(1/3)(1/3) = 3/9
P(7) = P(3 and 4 OR 4 and 3) = (1/3)(1/3)+(1/3)(1/3) = 2/9
P(8) = P(4 and 4) = (1/3)(1/3) = 1/9

Note that the sum of those probabilities is 9/9=1, giving us some reassurance that our method of calculating them is correct.